# Coarse spaces using extended generalized eigenproblems for heterogeneous Helmholtz problems

### v1

@@ 1 Optimized Restricted Additive Schwarz method
where $\mu$ and $\nu$ are two variable but positive and bounded coefficients, $f\in L^{2}(\Omega)$ and $g\in L^{2}(\partial\Omega)$ model the source terms, and $\mathbf{n}$ is the outward unit normal vector to the boundary $\partial\Omega$.
The model problem (1) takes the following variational form
@@
$\overline{\Omega}=\bigcup_{j}\overline{\Omega_{j}},\qquad\chi_{j}\geq 0,\ \chi_{j}|_{\partial\Omega_{j}\setminus\partial\Omega}=0,\qquad\sum_{j}E_{j}\chi_{j}u|_{\Omega_{j}}=u,\ \forall u\in H^{1}(\Omega),$ (5)
where $E_{j}$ is an extension-by-continuity operator. More precisely, we require that $E_{j}:H^{1}(\Omega_{j})\to H^{1}({\Omega})$ continuously and such that $\operatorname{supp}(E_{j}u_{j})\subset\widetilde{\Omega}_{j}$ for any $u_{j}\in H^{1}(\Omega_{j})$ and for some extended open subdomain $\widetilde{\Omega}_{j}$ such that ${\Omega}_{j}\subsetneq\widetilde{\Omega}_{j}\subset\Omega$.
@@
$u^{n+1}_{\mathrm{I}}:=u^{n}_{\mathrm{I}}+\sum_{j}E_{j}\chi_{j}v_{j}^{n+1},\qquad n\geq 0,$ (6)
@@
$b_{\Omega_{j}}(v_{j}^{n+1},v_{j}^{t})=l_{\widetilde{\Omega}_{j}}(E_{j}v_{j}^{t})-a_{\widetilde{\Omega}_{j}}(u^{n}_{\mathrm{I}}|_{\widetilde{\Omega}_{j}},E_{j}v_{j}^{t}),\qquad\forall v^{t}_{j}\in H^{1}(\Omega_{j}),$ (7)
@@
By similar arguments as for the model problem, the local problem (7) is well-posed.
One-level error analysis. For $n\geq 0$, the one-level error is defined as
$e^{n+1}_{\mathrm{I}}:=u-u^{n+1}_{\mathrm{I}}=e^{n}_{\mathrm{I}}-\sum_{j}E_{j}\chi_{j}v_{j}^{n+1}=\sum_{j}E_{j}\chi_{j}(e^{n}_{\mathrm{I}}|_{\Omega_{j}}-v_{j}^{n+1}).$ (9)
@@
$b_{\Omega_{j}}(e^{n}_{\mathrm{I}}|_{\Omega_{j}}-v_{j}^{n+1},u_{j}^{t})=b_{\Omega_{j}}(e^{n}_{\mathrm{I}}|_{\Omega_{j}},v_{j}^{t})-a_{\widetilde{\Omega}_{j}}(e^{n}_{\mathrm{I}}|_{\widetilde{\Omega}_{j}},E_{j}v_{j}^{t}),\ \forall v_{j}^{t}\in H^{1}(\Omega_{j}),$ (10)
@@
$e^{n+1}_{\mathrm{I}}=\sum_{j}E_{j}\chi_{j}R_{j}e^{n}_{\mathrm{I}}|_{\widetilde{\Omega}_{j}},$ (11)
@@
$b_{\Omega_{j}}(R_{j}\widetilde{v}_{j},v_{j}^{t})=b_{\Omega_{j}}(\widetilde{v}_{j}|_{\Omega_{j}},v_{j}^{t})-a_{\widetilde{\Omega}_{j}}(\widetilde{v}_{j},E_{j}v_{j}^{t}),\ \forall v_{j}^{t}\in H^{1}(\Omega_{j}).$ (12)
@@
$\|e^{n+1}_{\mathrm{I}}\|_{\Omega}^{2}\leq k_{0}\sum_{j}\|\chi_{j}R_{j}e^{n}_{\mathrm{I}}|_{\widetilde{\Omega}_{j}}\|_{\Omega_{j}}^{2}\leq k_{0}\xi\sum_{j}\|e^{n}_{\mathrm{I}}|_{\widetilde{\Omega}_{j}}\|_{\widetilde{\Omega}_{j}}^{2}\leq k_{0}k_{1}\xi\|e^{n}_{\mathrm{I}}\|_{\Omega}^{2},$ (14)
@@
$k_{0}:=\!\!\sup_{\begin{subarray}{c}(u_{j})_{j}\\ 0\neq u_{j}\in H^{1}_{0}(\Omega_{j})\end{subarray}}\!\!\frac{\|\sum_{j}E_{j}u_{j}\|_{\Omega}^{2}}{\sum_{j}\|u_{j}\|_{\Omega_{j}}^{2}},\qquad k_{1}:=\!\!\sup_{0\neq u\in H^{1}(\Omega)}\!\!\frac{\sum_{j}\|u|_{\widetilde{\Omega}_{j}}\|_{\widetilde{\Omega}_{j}}^{2}}{\|u\|_{\Omega}^{2}},$ (15)
@@
The estimate (14) shows that a sufficient condition to have convergence of the algorithm is that $k_{0}k_{1}\xi<1$, which is unlikely to be satisfied. The two constants $k_{0}$ and $k_{1}$ can be bounded uniformly of the parameters $\mu,\nu,\omega$ of the problem and only depend respectively on the maximum number of neighbors for a subdomain (see e.g. (Dolean2015, , Lem. 7.9)) and the maximum number of subdomain intersection multiplicity (see e.g. (Dolean2015, , Lem. 7.13)). For reasonable partitions, both quantities can be bounded uniformly of the number of subdomains. On the other hand, $\xi$ depends on the parameters of the problem. In the next section, we propose the construction of a coarse space and associated two-level domain decomposition algorithm that allows to control each $\xi_{j}$ independently, hence $\xi$.
@@ 2 Two-level ORAS method
$\begin{cases}c_{\Omega_{j}}(\chi_{j}v_{j},\chi_{j}v_{j}^{t})-b_{\Omega_{j}}(v_{j}^{t},\sigma_{j})=0,&\quad\forall v_{j}^{t}\in H^{1}(\Omega_{j}),\\ -b_{\Omega_{j}}(v_{j},\sigma_{j}^{t})+b_{\Omega_{j}}(\widetilde{u}_{j}|_{\Omega_{j}},\sigma_{j}^{t})-a_{\widetilde{\Omega}_{j}}(\widetilde{u}_{j},E_{j}\sigma_{j}^{t})=0,&\quad\forall\sigma_{j}^{t}\in H^{1}(\Omega_{j}),\\ b_{\Omega_{j}}(\widetilde{u}_{j}^{t}|_{\Omega_{j}},\sigma_{j})-a_{\widetilde{\Omega}_{j}}(\widetilde{u}_{j}^{t},E_{j}\sigma_{j})=\lambda_{j}c_{\widetilde{\Omega}_{j}}(\widetilde{u}_{j},\widetilde{u}_{j}^{t}),&\quad\forall\widetilde{u}_{j}^{t}\in H^{1}(\widetilde{\Omega}_{j}).\end{cases}$ (18)
@@
The result is based on a Caccioppoli inequality which we establish next (see also (Ma2023, , Lem. 2.7)). From (12) we have, for any $\widetilde{u}_{j}\in H^{1}(\widetilde{\Omega}_{j})$,
@@
$\quad+(\mu\chi_{j}\nabla R_{j}\widetilde{u}_{j},R_{j}\widetilde{u}_{j}\nabla\chi_{j})_{L^{2}(\Omega_{j})^{d}}-\omega^{2}(\nu R_{j}\widetilde{u}_{j},\chi_{j}^{2}R_{j}\widetilde{u}_{j})_{L^{2}(\Omega_{j})}=0.$
@@
$u^{n+1}_{\mathrm{II}}:=u^{n}_{\mathrm{II}}+\sum_{j}E_{j}\chi_{j}w_{j}^{n+1}+w_{0}^{n+1},\qquad n\geq 0,$ (22)
@@
$b_{\Omega_{j}}(w_{j}^{n+1},w_{j}^{t})=l_{\widetilde{\Omega}_{j}}(E_{j}w_{j}^{t})-a_{\widetilde{\Omega}_{j}}(u^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}},E_{j}w_{j}^{t}),\qquad\forall w^{t}_{j}\in H^{1}(\Omega_{j}),$ (23)
@@
$a_{\Omega}(w_{0}^{n+1},w_{0}^{t})=l_{\Omega}(w_{0}^{t})-a_{\Omega}(u^{n}_{\mathrm{II}}+\sum_{j}E_{j}\chi_{j}w_{j}^{n+1},w_{0}^{t}),\quad\forall w_{0}^{t}\in U_{\tau}.$ (24)
@@
Assumption 1 implies that the problem (24) is well-posed. The validity of Assumption 1 can be argued for large enough coarse spaces $U_{\tau}$ (i.e. taking the user-controlled parameter $\tau$ sufficiently small). For instance, well-posedness of the coarse problem and a quasi-optimality estimate implying (27) can be obtained using the so-called Aubin–Nitsche duality trick and Schatz argument, see (Galkowski2025, , Lemma 5.1).
Two-level error analysis. For $n\geq 0$, the two-level error is defined as
$e^{n+1}_{\mathrm{II}}:=u-u^{n+1}_{\mathrm{II}}=u-u^{n}_{\mathrm{II}}-\sum_{j}E_{j}\chi_{j}w_{j}^{n+1}-w_{0}^{n+1}.$ (28)
@@
$b_{\Omega_{j}}(e^{n}_{\mathrm{II}}|_{\Omega_{j}}-w_{j}^{n+1},w_{j}^{t})=b_{\Omega_{j}}(e^{n}_{\mathrm{II}}|_{\Omega_{j}},w_{j}^{t})-a_{\widetilde{\Omega}_{j}}(e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}},E_{j}w_{j}^{t}),\ \forall w_{j}^{t}\in H^{1}(\Omega_{j}),$ (29)
@@
$a_{\Omega}(w_{0}^{n+1},w_{0}^{t})=a_{\Omega}(\sum_{j}E_{j}\chi_{j}(e^{n}_{\mathrm{II}}|_{\Omega_{j}}-w_{j}^{n+1}),w_{0}^{t}),\quad\forall w_{0}^{t}\in U_{\tau},$ (30)
@@
$e^{n+1}_{\mathrm{II}}=(\mathrm{I}-P_{0})\sum_{j}E_{j}\chi_{j}R_{j}e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}},$ (31)
@@
We can now state our main result, which is the counterpart for the heterogeneous Helmholtz problem of the matrix formulation result stated in (Nataf2024, , Th. 2.12).
@@
$e^{n+1}_{\mathrm{II}}=(\mathrm{I}-P_{0})\sum_{j}E_{j}\chi_{j}R_{j}[\Pi_{j}+(\mathrm{I}-\Pi_{j})]e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}}=(\mathrm{I}-P_{0})\sum_{j}E_{j}\chi_{j}R_{j}(\mathrm{I}-\Pi_{j})e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}},$ (36)
and moreover (see e.g. (Dolean2015, , Lem. 7.7))
@@
The error can hence be estimated as follows
$\|e^{n+1}_{\mathrm{II}}\|_{\Omega}^{2}$ $\leq\sigma^{2}\Big\|\sum_{j}E_{j}\chi_{j}R_{j}(\mathrm{I}-\Pi_{j})e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}}\Big\|_{\Omega}^{2}\leq\sigma^{2}k_{0}\sum_{j}\|\chi_{j}R_{j}(\mathrm{I}-\Pi_{j})e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}}\|_{\Omega_{j}}^{2}$
$\leq\sigma^{2}k_{0}\tau\sum_{j}\|e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}}\|_{\widetilde{\Omega}_{j}}^{2}\leq\sigma^{2}k_{0}k_{1}\tau\|e^{n}_{\mathrm{II}}\|_{\Omega}^{2}.$ (38)
In practice, Krylov acceleration is used, and one iteration of the above algorithm (22) is used as a preconditioner. Theorem 2.1 also implies convergence of the Krylov iterative method.

### v2

@@ 1 Optimized Restricted Additive Schwarz method
where $\mu$ and $\nu$ are two variable but positive and bounded coefficients, $f\in L^{2}(\Omega)$ and $g\in L^{2}(\partial\Omega)$ model the source terms, and $\mathbf{n}$ is the outward unit normal vector to the boundary $\partial\Omega$. The model problem (1) takes the following variational form
@@
$\overline{\Omega}=\textstyle\bigcup_{j}\overline{\Omega_{j}},\qquad\chi_{j}\geq 0,\ \chi_{j}|_{\partial\Omega_{j}\setminus\partial\Omega}=0,\qquad\textstyle\sum_{j}E_{j}\chi_{j}u|_{\Omega_{j}}=u,\ \forall u\in H^{1}(\Omega),$ (5)
where $E_{j}$ is an extension-by-continuity operator. More precisely, we require that $E_{j}:H^{1}(\Omega_{j})\to H^{1}({\Omega})$ continuously, $E_{j}u_{j}|_{\Omega_{j}}=u_{j}$ and $\operatorname{supp}(E_{j}u_{j})\subset\widetilde{\Omega}_{j}$ for any $u_{j}\in H^{1}(\Omega_{j})$ and for some extended open subdomain $\widetilde{\Omega}_{j}$ such that ${\Omega}_{j}\subsetneq\widetilde{\Omega}_{j}\subset\Omega$.
@@
$u^{n+1}_{\mathrm{I}}:=u^{n}_{\mathrm{I}}+\textstyle\sum_{j}E_{j}\chi_{j}v_{j}^{n+1},\qquad n\geq 0,$ (6)
@@
$b_{\Omega_{j}}(v_{j}^{n+1},v_{j}^{t})=l_{\Omega}(E_{j}v_{j}^{t})-a_{\Omega}(u^{n}_{\mathrm{I}},E_{j}v_{j}^{t}),\qquad\forall v^{t}_{j}\in H^{1}(\Omega_{j}),$ (7)
@@
By similar arguments as for the model problem, the local problem (7) is well-posed, and we remark that the right-hand-side in (7) can be computed locally in $\widetilde{\Omega}_{j}$ so that $v_{j}^{n+1}$ is equivalently solution of
$b_{\Omega_{j}}(v_{j}^{n+1},v_{j}^{t})=l_{\widetilde{\Omega}_{j}}((E_{j}v_{j}^{t})|_{\widetilde{\Omega}_{j}})-a_{\widetilde{\Omega}_{j}}(u^{n}_{\mathrm{I}}|_{\widetilde{\Omega}_{j}},(E_{j}v_{j}^{t})|_{\widetilde{\Omega}_{j}}),\qquad\forall v^{t}_{j}\in H^{1}(\Omega_{j}).$ (9)
For general partitions (where at least three subdomains overlap in some part of the domain $\Omega$), the sequence of $u_{\mathrm{I}}^{n}$ depends a priori on the choice of $E_{j}$ (which is not unique), since it affects the right-hand-side of (7) (in (6), the presence of the partition-of-unity $\chi_{j}$ implies that $E_{j}$ extends by zero, which is uniquely defined).
One-level error analysis. For $n\geq 0$, the one-level error $e^{n}_{\mathrm{I}}:=u-u^{n}_{\mathrm{I}}$ satisfies
$e^{n+1}_{\mathrm{I}}=u-u^{n+1}_{\mathrm{I}}=e^{n}_{\mathrm{I}}-\textstyle\sum_{j}E_{j}\chi_{j}v_{j}^{n+1}=\textstyle\sum_{j}E_{j}\chi_{j}(e^{n}_{\mathrm{I}}|_{\Omega_{j}}-v_{j}^{n+1}).$ (10)
@@
$b_{\Omega_{j}}(e^{n}_{\mathrm{I}}|_{\Omega_{j}}-v_{j}^{n+1},v_{j}^{t})=b_{\Omega_{j}}(e^{n}_{\mathrm{I}}|_{\Omega_{j}},v_{j}^{t})-a_{\widetilde{\Omega}_{j}}(e^{n}_{\mathrm{I}}|_{\widetilde{\Omega}_{j}},(E_{j}v_{j}^{t})|_{\widetilde{\Omega}_{j}}),\ \forall v_{j}^{t}\in H^{1}(\Omega_{j}),$ (11)
@@
$e^{n+1}_{\mathrm{I}}=\textstyle\sum_{j}E_{j}\chi_{j}R_{j}e^{n}_{\mathrm{I}}|_{\widetilde{\Omega}_{j}},$ (12)
@@
$b_{\Omega_{j}}(R_{j}\widetilde{v}_{j},v_{j}^{t})=b_{\Omega_{j}}(\widetilde{v}_{j}|_{\Omega_{j}},v_{j}^{t})-a_{\widetilde{\Omega}_{j}}(\widetilde{v}_{j},(E_{j}v_{j}^{t})|_{\widetilde{\Omega}_{j}}),\qquad\forall v_{j}^{t}\in H^{1}(\Omega_{j}).$ (13)
@@
$\|e^{n+1}_{\mathrm{I}}\|_{\Omega}^{2}\leq k_{0}\textstyle\sum_{j}\big\|\chi_{j}R_{j}e^{n}_{\mathrm{I}}|_{\widetilde{\Omega}_{j}}\big\|_{\Omega_{j}}^{2}\leq k_{0}\xi\textstyle\sum_{j}\big\|e^{n}_{\mathrm{I}}|_{\widetilde{\Omega}_{j}}\big\|_{\widetilde{\Omega}_{j}}^{2}\leq k_{0}k_{1}\xi\|e^{n}_{\mathrm{I}}\|_{\Omega}^{2},$ (15)
@@
$k_{0}:=\!\!\sup_{\begin{subarray}{c}(u_{j})_{j}\\ 0\neq u_{j}\in H^{1}_{0}(\Omega_{j})\end{subarray}}\!\!\frac{\|\textstyle\sum_{j}E_{j}u_{j}\|_{\Omega}^{2}}{\textstyle\sum_{j}\|u_{j}\|_{\Omega_{j}}^{2}},\qquad k_{1}:=\!\!\sup_{0\neq u\in H^{1}(\Omega)}\!\!\frac{\textstyle\sum_{j}\|u|_{\widetilde{\Omega}_{j}}\|_{\widetilde{\Omega}_{j}}^{2}}{\|u\|_{\Omega}^{2}},$ (16)
@@
The estimate (15) shows that a sufficient condition to have convergence of the algorithm is that $k_{0}k_{1}\xi<1$, which is unlikely to be satisfied. The two constants $k_{0}$ and $k_{1}$ can be bounded uniformly of the parameters $\mu,\nu,\omega$ of the problem and only depend respectively on the maximum number of neighbors for a subdomain (see e.g. (Dolean2015, Lem. 7.9)) and the maximum number of subdomain intersection multiplicity (see e.g. (Dolean2015, Lem. 7.13)). For reasonable partitions, both quantities can be bounded uniformly of the number of subdomains. On the other hand, $\xi$ depends on the parameters of the problem. In the next section, we propose the construction of a coarse space and associated two-level domain decomposition algorithm that allows to control each $\xi_{j}$ independently, hence $\xi$.
@@ 2 Two-level ORAS method
$\begin{cases}c_{\Omega_{j}}(\chi_{j}v_{j},\chi_{j}v_{j}^{t})-b_{\Omega_{j}}(v_{j}^{t},\sigma_{j})=0,&\forall v_{j}^{t}\in H^{1}(\Omega_{j}),\\ -b_{\Omega_{j}}(v_{j},\sigma_{j}^{t})+b_{\Omega_{j}}(\widetilde{u}_{j}|_{\Omega_{j}},\sigma_{j}^{t})-a_{\widetilde{\Omega}_{j}}(\widetilde{u}_{j},(E_{j}\sigma_{j}^{t})|_{\widetilde{\Omega}_{j}})=0,&\forall\sigma_{j}^{t}\in H^{1}(\Omega_{j}),\\ b_{\Omega_{j}}(\widetilde{u}_{j}^{t}|_{\Omega_{j}},\sigma_{j})-a_{\widetilde{\Omega}_{j}}(\widetilde{u}_{j}^{t},(E_{j}\sigma_{j})|_{\widetilde{\Omega}_{j}})=\lambda_{j}c_{\widetilde{\Omega}_{j}}(\widetilde{u}_{j},\widetilde{u}_{j}^{t}),&\forall\widetilde{u}_{j}^{t}\in H^{1}(\widetilde{\Omega}_{j}).\end{cases}$ (19)
@@
The result is based on a Caccioppoli inequality which we establish next (see also (Ma2023, Lem. 2.7)). From (13) we have, for any $\widetilde{u}_{j}\in H^{1}(\widetilde{\Omega}_{j})$,
@@
$\quad+(\mu\chi_{j}\nabla R_{j}\widetilde{u}_{j},R_{j}\widetilde{u}_{j}\nabla\chi_{j})_{L^{2}(\Omega_{j})^{d}}-\omega^{2}(\nu R_{j}\widetilde{u}_{j},\chi_{j}^{2}R_{j}\widetilde{u}_{j})_{L^{2}(\Omega_{j})}=0.$ (21)
@@
$u^{n+1}_{\mathrm{II}}:=u^{n}_{\mathrm{II}}+\textstyle\sum_{j}E_{j}\chi_{j}w_{j}^{n+1}+w_{0}^{n+1},\qquad n\geq 0,$ (24)
@@
$b_{\Omega_{j}}(w_{j}^{n+1},w_{j}^{t})=l_{\Omega}(E_{j}w_{j}^{t})-a_{\Omega}(u^{n}_{\mathrm{II}},E_{j}w_{j}^{t}),\qquad\forall w^{t}_{j}\in H^{1}(\Omega_{j}),$ (25)
@@
$a_{\Omega}(w_{0}^{n+1},w_{0}^{t})=l_{\Omega}(w_{0}^{t})-a_{\Omega}(u^{n}_{\mathrm{II}}+\textstyle\sum_{j}E_{j}\chi_{j}w_{j}^{n+1},w_{0}^{t}),\quad\forall w_{0}^{t}\in U_{\tau}.$ (26)
@@
Assumption 1 implies that the problem (26) is well-posed. The validity of Assumption 1 can be argued for large enough coarse spaces $U_{\tau}$ (i.e. taking the user-controlled parameter $\tau$ sufficiently small). For instance, well-posedness of the coarse problem and a quasi-optimality estimate implying (29) can be obtained using the so-called Aubin–Nitsche duality trick and Schatz argument, see (Galkowski2025, Lemma 5.1).
Two-level error analysis. For $n\geq 0$, the two-level error $e^{n}_{\mathrm{II}}:=u-u^{n}_{\mathrm{II}}$ satisfies
$e^{n+1}_{\mathrm{II}}=u-u^{n+1}_{\mathrm{II}}=u-u^{n}_{\mathrm{II}}-\textstyle\sum_{j}E_{j}\chi_{j}w_{j}^{n+1}-w_{0}^{n+1}.$ (30)
@@
$b_{\Omega_{j}}(e^{n}_{\mathrm{II}}|_{\Omega_{j}}-w_{j}^{n+1},w_{j}^{t})=b_{\Omega_{j}}(e^{n}_{\mathrm{II}}|_{\Omega_{j}},w_{j}^{t})-a_{\widetilde{\Omega}_{j}}(e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}},(E_{j}w_{j}^{t})|_{\widetilde{\Omega}_{j}}),\ \forall w_{j}^{t}\in H^{1}(\Omega_{j}),$ (31)
@@
$a_{\Omega}(w_{0}^{n+1},w_{0}^{t})=a_{\Omega}(\textstyle\sum_{j}E_{j}\chi_{j}(e^{n}_{\mathrm{II}}|_{\Omega_{j}}-w_{j}^{n+1}),w_{0}^{t}),\quad\forall w_{0}^{t}\in U_{\tau},$ (32)
@@
$e^{n+1}_{\mathrm{II}}=(\mathrm{I}-P_{0})\textstyle\sum_{j}E_{j}\chi_{j}R_{j}e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}},$ (33)
@@
We can now state our main result, which is the counterpart for the heterogeneous Helmholtz problem of the matrix formulation result stated in (Nataf2025, Th. 2.12).
@@
$e^{n+1}_{\mathrm{II}}=(\mathrm{I}-P_{0})\textstyle\sum_{j}E_{j}\chi_{j}R_{j}[\Pi_{j}+(\mathrm{I}-\Pi_{j})]e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}}=(\mathrm{I}-P_{0})\textstyle\sum_{j}E_{j}\chi_{j}R_{j}(\mathrm{I}-\Pi_{j})e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}},$ (38)
and moreover (see e.g. (Dolean2015, Lem. 7.7))
@@
Recalling (16) and (35), the error can hence be estimated as follows
$\|e^{n+1}_{\mathrm{II}}\|_{\Omega}^{2}$ $=\big\|(\mathrm{I}-P_{0})\textstyle\sum_{j}E_{j}\chi_{j}R_{j}(\mathrm{I}-\Pi_{j})e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}}\big\|_{\Omega}^{2}$
$\leq\sigma^{2}\;\big\|\textstyle\sum_{j}E_{j}\chi_{j}R_{j}(\mathrm{I}-\Pi_{j})e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}}\big\|_{\Omega}^{2}\leq\sigma^{2}k_{0}\;\ \textstyle\sum_{j}\big\|\chi_{j}R_{j}(\mathrm{I}-\Pi_{j})e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}}\big\|_{\Omega_{j}}^{2}$
$\leq\sigma^{2}k_{0}\tau\;\textstyle\sum_{j}\big\|e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}}\big\|_{\widetilde{\Omega}_{j}}^{2}\leq\sigma^{2}k_{0}k_{1}\tau\;\|e^{n}_{\mathrm{II}}\|_{\Omega}^{2}.$ (40)
In practice, Krylov acceleration is used, and one iteration of the above algorithm (24) is used as a preconditioner. Theorem 2.1 also implies convergence of the Krylov iterative method with this preconditioner.